Grade 12th passAlgebra

Name the property under multiplication used in each of the following.
-4/5 × 1 = 1 × (-4/5) = -4/5

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

-4/5×1=1×(-4/5)=-4/5
Here 1 is the multiplicative identity.

​Thanks